### v1

# Bright single-photon source in a silicon chip by nanoscale positioning of a color center in a microcavity

Abstract
We present an all-silicon source of near-infrared linearly-polarized single photons, fabricated by nanoscale positioning of a color center in a silicon-on-insulator microcavity. The color center consists of a single W center, created at a well-defined position by Si+ ion implantation through a 150 nm-diameter nanohole in a mask. A circular Bragg grating cavity resonant with the W’s zero-phonon line at 1217 nm is fabricated at the same location as the nanohole. Under above-gap continuous-wave excitation, a very clean photon antibunching behavior () is observed over the entire power range, which highlights the absence of parasitic emitters. Purcell-enhancement of W’s zero-phonon emission provides both a record-high photoluminescence count rate among Si color centers (ca counts/s) and apparent Debye-Waller factor around 99%. We also demonstrate the triggered emission of single photons with 93% purity under weak pulsed laser excitation. At high pulsed laser power, we reveal a detrimental effect of repumping processes, that could be mitigated using selective pumping schemes in the future. These results represent a major step towards on-demand sources of indistinguishable near-infrared single photons within silicon photonics chips.
keywords:
[…]
Introduction
The generation, processing, and detection of single photons have enabled pioneering investigations in quantum physics, paving the way to groundbreaking technologies. Silicon quantum photonics, namely the processing of photonic quantum states in a silicon chip, is at the forefront of this revolution.1, 2 However, the potential of silicon photonics for large-scale quantum applications is limited by the difficulty to deterministically generate arbitrary large entangled photon states on-chip. A key requirement to overcome this limitation is to embed sources of indistinguishable single photons operating on demand. While integrated sources of entangled photon pairs based on spontaneous four-wave mixing have enabled major demonstrations,2 these sources suffer from an inherent interplay between efficiency and multi-photon pairs emission. Alternative approaches rely on fiber-coupling of an external source,3, 4 and hybrid integration of quantum emitters,5, 6, 7 albeit with limited efficiency due to coupling losses.
In this context, near-infrared color centers in silicon have emerged as a promising path towards deterministic single-photon sources (SPS) circumventing coupling losses. Due to their emission wavelength in the telecom bands, these point defects also appear as potential sources for chip-based quantum communications. Since the observation of photon antibunching for single color centers in 2020,8, 9 tremendous efforts have been devoted to optimizing their elaboration processes,10, 11, 12, 13, 14 and exploring their photophysical properties.15, 16, 17, 18, 19 In order to build a SPS, it is necessary to embed the emitter in a photonic cavity to achieve on-demand emission of single photons in a well-defined optical mode using the Purcell effect.20 Several single-mode sources in silicon have been demonstrated, based on a T center,21, 22 G center,23 G⋆ center,24 and erbium dopant.25 However, in spite of remarkable improvement of emission properties owing to a high Purcell factor, these sources suffer from significant limitations. The T center and erbium dopant are inherently slow emitters,21, 22 the G center has a poor quantum efficiency,26, 27, 23 and the fabrication process of the G⋆ center is not selective,15 resulting in poor purity due to the presence of parasitic defects.24 Moreover, in these first experiments the position of the emitter in the cavity is not controlled, hindering both the optimal exploitation of the Purcell effect and the reproducible fabrication of the SPS.
In recent work, Purcell enhancement has been demonstrated for an ensemble of W centers in circular Bragg grating (CBG) cavities in a silicon-on-insulator (SOI) film.28, 29 On the one hand, these results have positioned the W center as a promising candidate to build a SPS, owing to its high quantum efficiency () and relatively fast zero-phonon emission ( ns) in bulk silicon. 28 On the other hand, it confirms that CBG cavities, which have been widely used in other material systems,30, 31, 32, 33, 34, 35 are also attractive for building SPS on SOI chips.
In this letter, we present a SPS fabricated by nanoscale positioning of a single W center in a SOI CBG cavity. An array of W centers is created by Si+ ion implantation through a mask patterned with nanoholes,11 then an array of CBG cavities centered at the same location as the nanoholes and resonant with the W zero-phonon line (ZPL) is fabricated. The optical properties of a cavity-coupled W center are studied by photoluminescence (PL) spectroscopy and photon correlation measurements under above-gap continuous-wave (CW) and pulsed excitation. The purity of the source is analyzed as a function of the laser excitation power.
Sample design and fabrication
A 480 nm-thick mask of polymethyl methacrylate (PMMA) is deposited onto a commercial piece of SOI wafer. Arrays of nanoholes defined by electron beam lithography are etched into the mask The sample is then broad-beam implanted with Si+ ions (Figure 1(a)). A similar approach has already proven efficient for the fabrication of NV centers at preset positions in diamond,36, 37 and, more recently, G and W centers in silicon.11 The PMMA mask is then removed by stripping with acetone and the sample is annealed at 200 °C for 30 minutes in N2 atmosphere (Figure 1(b)). This second step leads to the activation of the W center,38, 39 whose microscopic structure (consisting of three interstitial Si atoms) is shown in the inset. A characterization of the emitters in unpatterned SOI is presented in Section 3 of the Supporting Information. CBG cavities centered at the same location as the nanoholes are fabricated by electron beam lithography and reactive ion etching (Figure 1(c)). This method is well-tried for the high-yield fabrication of SPS based on III-V semiconductor quantum dots (QD).40, 41 A map of the electromagnetic field in the central disk of the cavity is shown in the inset. After fabrication, the sample is first inspected by scanning electron microscopy, to confirm successful fabrication according to design parameters (see Figure S11 in the Supporting Information).
A sectional view of a CBG cavity containing a single W center is schematically represented in Figure 1(d), with the relevant dimensions. The red double-arrow represents the orientation of the transition dipole along the direction.16 It is well-known for CBG cavities that the Purcell enhancement factor , defined as the ratio between the emission rate in the cavity and bulk semiconductor, is heavily dependent on the position of the emitter.40 This sensivity was investigated for our emitter-cavity system using numerical finite-difference time-domain simulations (FDTD, see Section 4 in the Supporting Information). Figure 1(e) shows the Purcell enhancement factor for the ZPL as a function of the distance between the center of the cavity and the position of the dipole, in the coordinate system defined in Figure 1(d). A maximum enhancement factor of 13 is obtained at the center of the cavity. decreases rapidly against and reaches 0.3 for nm, emphasing the requirement for high positioning accuracy. Considering the simulated lateral spread of the ions implanted through a nanohole of 150 nm diameter, we estimate that the W centers have near-unity probability to be located less than 120 nm away from the center of the cavity, i.e. in the vicinity of the main antinode of the electromagnetic field (see Section 2 in the Supporting Information).
The collection efficiency for a cavity-coupled W center was also investigated using FDTD, considering the configuration of our measurement setup. The fraction of dipole emission radiated through a numerical aperture of 0.65 at normal incidence is plotted against in Figure 1(e). A collection efficiency of 51 is obtained at the center of the cavity. Note that this value could be increased to nearly with the addition of a high-reflectivity mirror below the cavity.42 The curves also reveal the robustness of the collection efficiency against spatial detuning (with for nm for each three coordinates), except for the "pit" around = 120 nm where the emitter is at an antinode of the electromagnetic field. This feature makes CBG cavities an attractive microstructure to easily increase the photon detection rate from quantum emitters in SOI. Similarly, the collection efficiency is also found robust against spectral detuning (see Figure S7 in the Supporting Information).
Experimental results
After fabrication, the sample is characterized by mapping the PL intensity in the emission range of W centers. On PL maps, CBG cavities that contain W centers appear as bright spots on a black background (Figure 2(a)). As the number of W centers formed through localized ion implantation is a random variable, we have thus to identify the CBG cavities that contain a single W center. Since the PL emission of single W centers, collected under normal incidence, is linearly polarized along either [110] or [1] crystal axes of Si, we identify very easily the CBG cavities that are the most likely to contain a single W emitter, by comparing two polarized-resolved PL maps, obtained for those crossed polarizations (see Section 5 in the Supporting Information). Photon-autocorrelation experiments in the Hanbury Brown and Twiss configuration are then performed to confirm the presence of one and only one W center in the cavity.
We display in Figure 2 a typical set of results obtained for a CBG cavity containing a single W center under above-gap CW excitation (see Section 1 in the Supporting Information for details about the measurement configuration). A measurement of the polarization diagram of the PL evidences linear polarization along the [110] axis of the silicon crystal with 97.2 0.7 visibility (see Figure S9 in the Supporting Information). We observe of a clean antibunching dip in the photon-autocorrelation histogram shown in Figure 2(c, d). It is worth emphasizing that this low value of is obtained without any correction for detectors dark counts or background emission. A bunching behaviour () is observed at non-zero delay. This phenomenon, already reported for single W centers,16, 11 is attributed to a metastable state coupled to the excited state of the emitter.
In order to highlight cavity effects on the single W center emission, we also show in Figure 2(b) a study of the PL intensity as a function of the CW pump power. We observe a well defined saturation behaviour and a count rate at saturation around Mcps. This value is remarkably high compared to previously published results for other color centers in SOI, as shown in Table 1. Importantly, we observe a very low over the full range of pump powers up to the saturation plateau ( is equal to 0.03 at 3 and to 0.06 at 9 ), unlike most previous reports. This important observation confirms that our fabrication process does not produce parasitic emitters, besides the desired isolated W center. The beneficial effect of the CBG cavity is most obviously evidenced by comparing the latter PL count rate at saturation to the one of isolated W centers in unpatterned SOI which is typically around 3.2 0.8 kcps after correction for background counts rate (see Section 3 in the Supporting Information). This represents an enhancement factor of around 400 for the PL count rate thanks to cavity effects. As shown below, this increase results from the combination of Purcell-enhancement of zero-photon emission into the CBG cavity mode and the small divergence of the radiation pattern of this mode in the far-field, providing efficient collection of zero-phonon emission.
| Em. | Cav. | (kcps) | (ns) | Pos. | Reference | |
| G⋆ | 2D PhC | 20 | 6.7 | 0.30 | No | Redjem et al.24 |
| G | 2D PhC | 20 | 6 | 0.03 | No | Saggio et al. see SI in ref.[23] |
| T | 1D PhC | 20 | 169 | 0.07 | No | Islam et al.21 |
| 236 | 0.35 | |||||
| W | CBG | 940 | 7 | 0.03 | Yes | This work |
| 1200 | 0.06 |
To understand quantitatively cavity effects, we investigated the properties of the cavity mode and the spectral resonance of the emitter-cavity system. The modal properties have been probed by performing reflection spectroscopy under white light illumination. As shown in Figure 3(b), the reflectivity curve features a dip around 1218 nm, fitting to a Lorentzian profile with a full width at half maximum (FWHM) nm, corresponding to a quality factor . This value is in good agreement with the quality factor estimate of 167 obtained from FDTD simulations. The CBG cavity mode is therefore nearly perfectly matched to the ZPL emission of zero-phonon emission line of W centers in unstrained silicon,16 as confirmed from a comparison of reflectivity and PL emission spectra in Figure 3(a). Noticeably, this PL spectrum over the 1140-1450 nm spectral range reveals that zero-phonon emission accounts for of the PL intensity, value to be compared with 39 for isolated W in unpatterned SOI, which unveils a selective enhancement of ZPL emission thanks to the Purcell effect.
We study now the recombination dynamics of the W center to estimate the magnitude of the Purcell effect. The PL decay of the emitter under pulsed excitation at 17.5 W with a 5 MHz repetition rate is displayed in Figure 3(c). The intensity decays nearly mono-exponentially with a characteristic time ns. This value should be compared with the characteristic time of around 34 ns obtained for ensembles of W centers in unpatterned SOI.39, 28, 29 The acceleration of the decay compared to this reference value has two possible origins: Purcell enhancement of zero-phonon emission, and/or activation of extra non-radiative recombination channels due to the interaction between the emitter and its environment. This second possibility cannot be neglected a priori as a significant lifetime dispersion has been reported for single W centers in 60 nm-thick SOI,16 although this dispersion can likely be attributed to surface effects in such a thin SOI layer. A detailed analysis of the photon budget, from emission by W centers to detection by our single photon detectors, provides a lower-bound estimate around 1/18 ns-1 for the zero-phonon emission rate (see Section 6 in the Supporting Information). By comparison with a previous estimate of 1/134 ns-1 in bulk Si,28 we conclude that the zero-phonon emission rate is increased by a factor . Therefore, the acceleration of the PL decay results predominantly (and possibly entirely) from the Purcell effect. Additionally, according to Figure 1(e), we can also conclude that the emitter is positionned less than 100 nm away from the cavity center.
The bright emission of antibunched photons under CW excitation unveils the strong potential of W centers for building a deterministic SPS in SOI. Along this line, we have explored the triggered emission of single photons under pulsed excitation. Figure 4(a) shows uncorrected autocorrelation histograms obtained for different pump powers for a 5MHz repetition rate. In this figure, histograms are normalized such that the area of all peaks would be equal to 1 for a coherent source. The normalized area as a function of time delay , where is the laser repetition period, of each peak is given by the coordinate of the squares on the right axis. In the low power regime, the peak at zero-delay has an area , demonstrating the triggered emission of single-photon pulses with 93 purity. As for the measurement under CW excitation, a bunching behaviour is observed (>1), indicating photoblinking of the source. Note that this effect is a rather common issue for solid-state quantum emitters and is expected for W, as this color center possesses a metastable excited state.16 The area of the peaks follows an exponential law given by:
where and are the average times of stay of the emitter in the "bright" and "dark" states, respectively.43 For = 0.5, 4.4, and 17.5 W, = 2290, 341, and 153 ns, and = 13200, 2200, and 741 ns, pointing towards an acceleration of the blinking with increasing power. Interestingly, the peaks for have a smaller area than expected from the model in the low-power regime. This effect, which has been reported for single InAs QD under above-gap excitation,44 may be related to a two-step transition from the ground state to the excited state, which is indeed predicted by ab-initio calculation.16
The area of the central peak increases with power, as shown in Figure 4(b). For = 17.5 W, the area of the peak reaches 0.51, meaning that about two photons are emitted on average per excitation pulse. A zoom on the peak with 20 ps time bins (Figure 4(c)) reveals an antibunching dip with a characteristic time of 60 ps (that disappears in Figure 4(b) due to rebinning). This antibunching behaviour indicates that the two photons are emitted successively. An explanation is that, under strong above-gap pumping condition, a second exciton can be captured after the emission of a first photon, as reported for other quantum emitters.45 This phenomenon is not surprizing here, due to the optical injection of a large population of carriers by the pump pulse in the central disk of the cavity. As for QD-SPS, we expect to overcome this limitation through the implementation of resonant or quasi-resonant pumping schemes.
Present limitations and development prospects
Under pulsed excitation at 5MHz frequency, the PL count rate reaches 52 kcps at saturation. This value corresponds, after taking into account transmission losses, to a source efficiency (as measured at the first lens of our collection optics) comprised between 0.03 and 0.06 photon per pulse (see Section 6 of the SI). The main factors limiting at this stage are easily identified: (1) W is in its bright state one sixth of the time, which reduces its average efficiency by the same amount; (2) our CBG cavity funnels only one-half of the emission towards our collection optics; (3) parasitic recombination channels (non-radiative ones and to a lesser extent phonon-assisted recombination) reduce the quantum efficiency of the zero-phonon emission to a value comprised between 0.36 and 0.68.
While point (1) and the issue of blinking has already been discussed, we see that points (2) and (3) can be radically improved by replacing our low-Fp CBG cavity by a high-Fp waveguide-coupled photonic crystal cavity, such as a nanobeam cavity. 46, 21, 22 Proper engineering of light extraction towards the waveguide mode and a large but realistic Purcell factor around 150 would bring the zero-phonon emission rate above 109 s-1. This would ensure that the quantum efficiency of the ZPL, the spontaneous emission factor and the efficiency of the waveguide-coupled SPS are all near unity. Single photon coherence and indistinguishability are also required properties for applications to quantum simulations and computing. Until now, Hong-Ou-Mandel interference have only been demonstrated over relatively narrow time-windows, shorter than the emitter lifetime. 47 Combining such a large Purcell acceleration of zero-phonon emission, with strategies aiming at reducing dephasing processes (such as quasi-resonant pumping or the application of a static electric field), will likely permit to generate highly-indistinguishable photons on demand.
From a technological point of view, our strategy for positioning single W centers is fully compatible with the geometry of photonic crystal cavities. In order to fully benefit from their large Purcell-enhancement factor, one has to implement means enabling a fine control of the spectral detuning between W’s ZPL and the high-Q cavity mode. In this context, electrical and strain tuning of the ZPL have been recently demonstrated for G centers. 48, 49
Conclusion
In conclusion, we have demonstrated a source of near-infrared linearly-polarized single photons obtained by nanoscale positioning of a W center in a SOI CBG cavity. Si+ implantation through a mask patterned with nanoholes allows for selective creation of single W centers, positioned in the vicinity of the main antinode of the resonant mode. The cavity-coupled emitter has a PL intensity at saturation of around 1.2 Mcps together with a value of around 0.06 under above-gap CW excitation, with a ZPL fraction of 98.6. Under weak pumping condition, the triggered emission of single-photon pulses is demonstrated with 93 purity. These results highlight the great potential of W centers for silicon quantum photonics. They pave the way towards the scalable integration of bright sources of indistinguishable single photons on integrated photonic chips, based on single W centers deterministically positioned in high-Fp waveguide-coupled photonic crystal nanocavities.
Methods
Sample preparation.
The silicon layer and buried oxide layer of our SOI have a thickness of 220 nm and 2 m, respectively.
The PMMA layer is chosen thick enough (480 nm) to stop the Si+ ions with nearly 100 probability during the implantation (see Figure S3 in the Supporting Information).
The nanoholes etched into the mask have diameters ranging from 125 nm to 2 m (see Table S1 in the Supporting Information), and 10 m pitch between the holes.
The implantation of Si+ ions is performed at an energy of 70 keV, with a dose of 1011 cm-2 and 7° tilt to avoid channeling.
Experimental setup.
The full optical measurement setup is schematically represented in SI Section 1.
The sample is cooled at 10 K in a helium-flow cryostat (Oxford Instruments Microstat HiRes).
Continuous-wave and pulsed optical excitation are provided by a semiconductor laser (PicoQuant LDH-IB-485-B) operating around 485 nm.
In pulsed operation mode, the pulses have a full-width at half-maximum (FWHM) of 100 ps and tunable repetition rate.
The laser beam is coupled into a single-mode fiber for spatial filtering, and collimated using a microscope objective mounted onto a translation stage in order to finely control the divergence of the beam.
The laser is focused onto the sample to a 1/e2-spot diameter of 700 nm using a microscope objective with a numerical aperture of 0.65 (Olympus LCPLN50XIR).
The objective is mounted onto a tri-axial positioner allowing for 50 nm translation steps for photoluminescence (PL) mapping.
A visualization line is accessible using a removable beamsplitter, in order to obtain an image of the sample on a CCD camera for alignment procedures.
The PL is collected using the microscope objective and separated from the laser using a dichroic mirror with a cutoff wavelength of 650 nm. The 1200 - 1300 nm spectral range is selected through a combination of longpass and shortpass filters, and the PL is focused into a single-mode fiber using a reflective collimator (Thorlabs RC04APC-P01). The PL is detected using superconducting nanowire single-photon detectors (SNSPD, ID Quantique ID281), with a detection efficiency of 81 for Detector 1 at 1310 nm (resp. 86 for Detector 2), dark count rate of 9 cps (resp. 8 cps), and timing jitter of 29 ps (resp. 30 ps). The polarization of the photons impinging on the SNSPD is controlled using a half-waveplate and a fiber polarization controller for optimized detection efficiency for both detectors. The photon detection events are counted by a time-correlated single-photon counting card (PicoQuant MultiHarp 150). For second-order correlation measurements, the card is operated in a time-tagged time-resolved mode.
For spectral measurements, the PL is dispersed by a grating with 900 lines/mm in a spectrometer (Horiba JY iHR320) and detected by a linear InGaAs array (Horiba 81043 Spectrum One). The system has a spectral resolution of 0.4 nm.
For polarization-resolved measurements, a motorized half-waveplate is placed in front of a polarizing beamsplitter cube (PBS). The PBS has a transmission of 98.2 for P-polarized light and 0.135 for S-polarized light at 1218 nm. The intensity is recorded as a function of the angle between the axis of the half-waveplate and the transmission axis of the PBS. The intensity is plotted as a function of 2 to reconstruct the polarization diagram of the PL.
For reflectometry measurements, white-light from a tungsten source is injected into a single-mode fiber and collimated using a reflective collimator. A 50:50 removable beamsplitter cube is used to inject the white beam into the excitation line, and the reflected light is collected using the same collection line as for the PL.
We thank Guillaume Cassabois and Vincent Jacques for fruitful discussions, as well as Krithika V. R. and Alicia Morandini for experimental support. This work is supported by the Plan France 2030 through the project OQuLuS (No. ANR-22-PETQ-0013), the French National Research Agency (ANR) through the projects OCTOPUS (No. ANR-18-CE47-0013-01) and WOUAH (No. ANR-24-CE47-4667), CEA through the PTC-MP “W-TeQ” internal project, and the European Research Council (ERC) under the European Union’s Horizon 2020 research and innovation programme (project SILEQS, Grant No. 101042075). B.L. is supported by the “Program QuantForm-UGA No. ANR-21-CMAQ-003 France 2030” and “Laboratoire d’Excellence LANEF No. ANR-10-LABX-51-01”. C.E. is supported by the MSCA Cofund QuanG (Grant Number: 101081458), funded by the European Union. The authors acknowledge the assistance of the staff of the Grenoble Advanced Technological Platform (PTA).
See the Supporting Information for additional details about simulation of ion implantation, optical properties of single W centers in unpatterned SOI, finite-difference time-domain simulations, single W centers hunting strategy, and estimation of the zero-phonon emission rate for the CBG-coupled W center.